OpenStudy (anonymous):

y varies inversely with twice the value of x. When y = 10, x = 3. Find y when x = 5. Choose one answer. a. 4 b. 6 c. 10 d. 12

myininaya (myininaya):

y=k/(2x) 10=k/(2*3) 6*10=k 60=k y=60/(2x)=30/x y=30/5=6

OpenStudy (anonymous):

y=k/(2x) 10=k/6 k=60 y=60/(2*5)=6 choice b

OpenStudy (anonymous):

x went up by 5/3, so y will go up by 3/5. 3/5 * 10 = 6